32/4+16-4*3=4
This problem must be solved using only 1 set of parenthesis.
Everyone in my family has tried and we can't get it.

Found 2 solutions by scott8148, nerdybill:
Answer by scott8148(5882)

everything after the fraction line on the left side goes in parentheses
___ remember PEMDAS

32/(4+16-4*3) = 4

32/(4+16-12) = 4

32/8 = 4

Answer by nerdybill(5404)

It would be:
32/(4+16-4*3)=4
.
Because then the order of precedence says multiplication first inside the parenthesis:
32/(4+16-12)=4
32/(20-12)=4
32/8=4
4=4